# thanks: M.C. and D.P. contributed equally to this work

### v3

@@ III.1 Greybody factors
$-\frac{h}{g}\partial^{2}_{t}\Upsilon_{s}+s\sqrt{\frac{f}{g}}\left(\frac{hg^{\prime}}{g}-h^{\prime}\right)\partial_{t}\Upsilon_{s}$
$+fh\partial_{r}^{2}\Upsilon_{s}+\left(\frac{hf^{\prime}}{2}+(s+1/2)\frac{fhg^{\prime}}{g}+(s+1)fh^{\prime}\right)\partial_{r}\Upsilon_{s}$
@@
$\left(s\frac{hfg^{\prime\prime}}{g}+\frac{3s-2s^{2}}{4}(2fh^{\prime\prime}+f^{\prime}h^{\prime})+\frac{s}{2}(\frac{hf^{\prime}g^{\prime}}{g}-\frac{fhg^{\prime 2}}{g^{2}})\right.$
$\left.+\frac{2s^{2}-s}{4}\frac{fh^{\prime 2}}{h}+\frac{2s^{2}+5s}{4}\frac{fg^{\prime}h^{\prime}}{g}\right)\Upsilon_{s}=0\,,$ (12)
@@
$A_{s}=\sqrt{\frac{f}{g}}\frac{1}{(gh)^{s}}\,,$ (14)
$B_{s}=\sqrt{fg}(gh)^{s}h\,,$ (15)
$C_{s}$ $=s\frac{fhg^{\prime\prime}}{g}+\frac{s}{2}\left(\frac{hf^{\prime}g^{\prime}}{g}-\frac{fhg^{\prime 2}}{g^{2}}\right)$
$+\frac{s(3-2s)}{4}\left(2fh^{\prime\prime}+f^{\prime}h^{\prime}\right)+\frac{s(2s-1)}{4}\frac{fh^{\prime 2}}{h}\,.$ (16)
With these definitions, the decoupled radial Teukolsky equation reduces to the following general form Arbey et al. (2021b):
$A_{s}(B_{s}R^{\prime}_{s})^{\prime}+\left[\frac{h}{g}\omega^{2}+i\omega s\sqrt{\frac{f}{g}}\left(h^{\prime}-\frac{hg^{\prime}}{g}\right)+C_{s}\right]R_{s}=0\,.$
@@
$\mathcal{A}_{s}=f^{2}h\,,$ (23)
$\mathcal{B}_{s}=\left(\left(s+\frac{1}{2}\right)\frac{f^{2}h\dot{g}}{g}+\frac{hf}{2}+(1+s)f^{2}\dot{h}\right)\,,$ (24)
$\mathcal{C}_{s}$ $=\frac{f}{4}\left(s(2s-1)\frac{f{\dot{h}}^{2}}{h}+2s(3-2s)f\ddot{h}-2s\frac{fh{\dot{g}}^{2}}{g^{2}}\right.$
$\left.+\frac{1}{g}\left(s(5+2s)f\dot{g}\dot{h}+2h\left(2\omega^{2}+2sf\ddot{g}\right.\right.\right.$ (25)
$\left.\left.\left.-s\dot{g}\left(2i\omega\sqrt{\frac{f}{g}+\dot{f}}\right)\right)\right)+sh\left((3-2s)\dot{f}+4i\omega\sqrt{\frac{f}{g}}\right)\right)\,,$
where the dot denotes a derivative with respect to the rescaled coordinate $x$. For completeness, we note that $f$ and $g$ as as function of $x$ for the metrics considered here are given by the following:
$\begin{split}&h_{\text{SV}}(x)=h_{\text{PK}}(x)=h_{\text{DR}}(x)=(x+1)^{2}+\ell^{2}\,,\\ &g_{\text{SV}}(x)=f_{SV}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &g_{\text{PK}}(x)=f_{\text{PK}}(x)=\frac{x}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &g_{\text{DR}}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &f_{\text{DR}}(x)=g_{\text{DR}}(x)\left(1+\frac{\ell}{\sqrt{\ell^{2}+(x+1)^{2}}}\right)^{-1}\,.\end{split}$ (26)
@@ Appendix A Asymptotic solutions to the radial Teukolsky equation
$\left[\omega^{2}+i\omega s\sqrt{\frac{f}{g}}\left(h^{\prime}-\frac{hg^{\prime}}{g}\right)\frac{g}{h}+C_{s}\frac{g}{h}-\frac{\sqrt{\beta},_{\star\star}}{\sqrt{\beta}}\right]Y_{s}=0\,,$
where $\beta\equiv(hg^{2})h=B_{s}/\sqrt{fg}$, and $,_{\star}$ denotes differentiation with respect to the tortoise coordinate $r_{\star}$. We now consider the $r\to+\infty$ and $r\to r_{H}$ limits separately.
@@ A.2 Asymptotic behaviour near the horizon
$Y_{s},_{\star\star}+\left[\left(\omega-\frac{is\sqrt{\frac{f}{g}}g^{\prime}}{4}\right)^{2}+\frac{g^{\prime}s}{4}\left(f^{\prime}-\frac{fg^{\prime}}{g}\right)\right]Y_{s}=0\,,$
@@
$Y_{s},_{\star\star}+\left(\omega-\frac{is\sqrt{\frac{f}{g}}g^{\prime}}{4}\right)^{2}Y_{s}=0\,,$ (44)
@@
$Y_{s}\sim\exp\left[\pm i\left(\omega-\frac{isg^{\prime}}{2}\sqrt{\frac{f}{g}}\right)r_{\star}\right]\,.$ (45)
@@
$Y_{s}\sim e^{\pm i\omega r_{\star}}\exp\left[\pm\frac{sg^{\prime}}{2}\sqrt{\frac{f}{g}}\text{K}\ln(r-r_{H})\right]\,.$ (47)
@@ A.2.3 D’Ambrosio-Rovelli
with $r_{H}^{\text{DR}}=\sqrt{4M^{2}-\ell^{2}}$. From Eq. (59) it follows that $\text{K}=2\sqrt{2}M^{3/2}(2M-\ell)$, which cancels once again with the $g^{\prime}\sqrt{f/g}$ term in Eq. (47). Analogously to the previous case, one can therefore conclude that the asymptotic solutions to the radial Teukolsky equation for $R_{s}$ in the near horizon region are indeed given by Eq. (55), thereby confirming the asymptotic scaling quoted in Eq. (20).

### v4

@@ III.1 Greybody factors
$-\frac{h}{f}\partial^{2}_{t}\Upsilon_{s}+s\sqrt{\frac{g}{f}}\left(\frac{hf^{\prime}}{f}-h^{\prime}\right)\partial_{t}\Upsilon_{s}$
$+gh\partial_{r}^{2}\Upsilon_{s}+\left(\frac{hg^{\prime}}{2}+(s+1/2)\frac{ghf^{\prime}}{f}+(s+1)gh^{\prime}\right)\partial_{r}\Upsilon_{s}$
@@
$\left(s\frac{hgf^{\prime\prime}}{f}+\frac{3s-2s^{2}}{4}(2gh^{\prime\prime}+g^{\prime}h^{\prime})+\frac{s}{2}(\frac{hg^{\prime}f^{\prime}}{f}-\frac{ghf^{\prime 2}}{f^{2}})\right.$
$\left.+\frac{2s^{2}-s}{4}\frac{gh^{\prime 2}}{h}+\frac{2s^{2}+5s}{4}\frac{gf^{\prime}h^{\prime}}{f}\right)\Upsilon_{s}=0\,,$ (12)
@@
$A_{s}=\sqrt{\frac{g}{f}}\frac{1}{(fh)^{s}}\,,$ (14)
$B_{s}=\sqrt{fg}(fh)^{s}h\,,$ (15)
$C_{s}$ $=s\frac{ghf^{\prime\prime}}{f}+\frac{s}{2}\left(\frac{hg^{\prime}f^{\prime}}{f}-\frac{ghf^{\prime 2}}{f^{2}}\right)$
$+\frac{s(3-2s)}{4}\left(2gh^{\prime\prime}+g^{\prime}h^{\prime}\right)+\frac{s(2s-1)}{4}\frac{gh^{\prime 2}}{h}$
$+\frac{s(2s+5)}{4}\frac{gf^{\prime}h^{\prime}}{f}-\lambda^{s}_{l}-2s\,.$ (16)
[…]
$A_{s}(B_{s}R^{\prime}_{s})^{\prime}+\left[\frac{h}{f}\omega^{2}+i\omega s\sqrt{\frac{g}{f}}\left(h^{\prime}-\frac{hf^{\prime}}{f}\right)+C_{s}\right]R_{s}=0\,.$
@@
$\mathcal{A}_{s}=g^{2}h\,,$ (23)
$\mathcal{B}_{s}=\left(\left(s+\frac{1}{2}\right)\frac{g^{2}h\dot{f}}{f}+\frac{hg\dot{g}}{2}+(1+s)g^{2}\dot{h}\right)\,,$ (24)
$\mathcal{C}_{s}$ $=\frac{g}{4}\left(s(2s-1)\frac{g{\dot{h}}^{2}}{h}+2s(3-2s)g\ddot{h}-2s\frac{gh{\dot{f}}^{2}}{f^{2}}\right.$
$\left.+\frac{1}{f}\left(s(5+2s)g\dot{f}\dot{h}+2h\left(2\omega^{2}+2sg\ddot{f}\right.\right.\right.$ (25)
$\left.\left.\left.-s\dot{f}\left(2i\omega\sqrt{\frac{g}{f}+\dot{g}}\right)\right)\right)+sh\left((3-2s)\dot{g}+4i\omega\sqrt{\frac{g}{f}}\right)-4\nu^{s}_{l}\right)\,,$
where the dot denotes a derivative with respect to the rescaled coordinate $x$ and $\nu^{s}_{l}=l(l+1)-s(s-1)$. For completeness, we note that $f$ and $g$ as as function of $x$ for the metrics considered here are given by the following:
$\begin{split}&h_{\text{SV}}(x)=h_{\text{PK}}(x)=h_{\text{DR}}(x)=(x+1)^{2}+\ell^{2}\,,\\ &g_{\text{SV}}(x)=f_{SV}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &g_{\text{PK}}(x)=f_{\text{PK}}(x)=\frac{x}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &f_{\text{DR}}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &g_{\text{DR}}(x)=f_{\text{DR}}(x)\left(1+\frac{\ell}{\sqrt{\ell^{2}+(x+1)^{2}}}\right)^{-1}\,.\end{split}$ (26)
@@ Appendix A Asymptotic solutions to the radial Teukolsky equation
$\left[\omega^{2}+i\omega s\sqrt{\frac{g}{f}}\left(h^{\prime}-\frac{hf^{\prime}}{f}\right)\frac{f}{h}+C_{s}\frac{f}{h}-\frac{\sqrt{\beta},_{\star\star}}{\sqrt{\beta}}\right]Y_{s}=0\,,$
where $\beta\equiv(hf^{2})h=B_{s}/\sqrt{fg}$, and $,_{\star}$ denotes differentiation with respect to the tortoise coordinate $r_{\star}$. We now consider the $r\to+\infty$ and $r\to r_{H}$ limits separately.
@@ A.2 Asymptotic behaviour near the horizon
$Y_{s},_{\star\star}+\left[\left(\omega-\frac{is\sqrt{\frac{g}{f}}f^{\prime}}{4}\right)^{2}+\frac{f^{\prime}s}{4}\left(g^{\prime}-\frac{gf^{\prime}}{f}\right)\right]Y_{s}=0\,,$
@@
$Y_{s},_{\star\star}+\left(\omega-\frac{is\sqrt{\frac{g}{f}}f^{\prime}}{4}\right)^{2}Y_{s}=0\,,$ (44)
@@
$Y_{s}\sim\exp\left[\pm i\left(\omega-\frac{isf^{\prime}}{2}\sqrt{\frac{g}{f}}\right)r_{\star}\right]\,.$ (45)
@@
$Y_{s}\sim e^{\pm i\omega r_{\star}}\exp\left[\pm\frac{sf^{\prime}}{2}\sqrt{\frac{g}{f}}\text{K}\ln(r-r_{H})\right]\,.$ (47)
@@ A.2.3 D’Ambrosio-Rovelli
with $r_{H}^{\text{DR}}=\sqrt{4M^{2}-\ell^{2}}$. From Eq. (59) it follows that $\text{K}=2\sqrt{2}M^{3/2}(2M-\ell)$, which cancels once again with the $f^{\prime}\sqrt{g/f}$ term in Eq. (47). Analogously to the previous case, one can therefore conclude that the asymptotic solutions to the radial Teukolsky equation for $R_{s}$ in the near horizon region are indeed given by Eq. (55), thereby confirming the asymptotic scaling quoted in Eq. (20).